OpenStudy (ranya99):

. A .22 rifle bullet, traveling at 350 m/s, strikes a block of soft wood, which it penetrates to a depth of 0.130 m. The block of wood is clamped in place and doesn’t move. The mass of the bullet is 1.80 g. Assume a constant retarding force. How much time is required for the bullet to stop? What force, in Newtons, does the wood exert on the bullet? 1kg = 1000g

rvc (rvc):

@Michele_Laino will help u with this m very sleepy atm

OpenStudy (ranya99):

@rvc alright thanks tho

OpenStudy (michele_laino):

here the work \(W\) by such retarding force \(F\) has to be equal to the kinetic energy \(T\) of the bullet, so we can write: \[\huge T = \frac{1}{2}m{v^2} = Fd\] where \(m\) is the mass of the bullet, and \(v\) is its speed

rvc (rvc):

thanks @Michele_Laino

OpenStudy (michele_laino):

of course, \(d\) is the depth

OpenStudy (ranya99):

so it would be like t=1/2(1.80)*(350)^2= F*0.130 right?

OpenStudy (michele_laino):

more precisely, we have to measure the bullet mass using Kg, so we can write: \[\Large \frac{1}{2} \cdot 0.0018 \cdot {350^2} = F \cdot 0.130\]

OpenStudy (michele_laino):

I got this: \[\large T = 110.25\;Joules\]

OpenStudy (michele_laino):

and: \[\Large F = \frac{{0.0018 \cdot {{350}^2}}}{{2 \cdot 0.13}} = ...?\]

OpenStudy (ranya99):

14.3?

OpenStudy (michele_laino):

I got this: \[\Large F = \frac{{0.0018 \cdot {{350}^2}}}{{2 \cdot 0.13}} = 848\;newtons\]

OpenStudy (ranya99):

ohh ok got it thanks

OpenStudy (michele_laino):

finally the requested time, in order to stop the bullet, is fiven by the subsequent computation: \[\Large \tau = \frac{{mv}}{F} = \frac{{0.0018 \cdot 350}}{{848}} = ...?\]

OpenStudy (michele_laino):

oops.. is given*

OpenStudy (ranya99):

0.000742

OpenStudy (michele_laino):

that's right!

OpenStudy (michele_laino):

it is \(\tau=0.000743 \) seconds

OpenStudy (ranya99):

okay, thank you so much for your help